# Universal Approximation Theory: The basic theory for deep learning-based computer vision models

### v2

@@ Abstract
Computer vision (CV) is one of the most crucial fields in artificial intelligence. In recent years, a variety of deep learning models based on convolution and Transformer have been designed to tackle diverse problems in CV. These algorithms have found practical applications in areas such as robotics and facial recognition. Despite the increasing power of current CV models, several fundamental questions remain unresolved: Why do convolutional neural networks (CNNs) require deep layers? What ensures the generalization ability of CNNs? Why do residual-based networks outperform fully convolutional networks like VGG? What is the fundamental difference between residual-based CNNs and Transformer-based networks? What are the differences between human vision and deep learning-based CVs? The root cause of these questions lies in the lack of a robust theoretical foundation for deep learning models in CV. To address these critical issues and techniques, we employ the Universal Approximation Theorem (UAT) to provide a theoretical basis for convolution- and Transformer-based models in CV. By doing so, we aim to elucidate these questions from a theoretical perspective.
@@ 1 Introduction
As a core branch of artificial intelligence, CV has a wide range of applications, encompassing tasks such as image segmentation (Minaee et al. 2021; Wang et al. 2022a), classification (Rawat and Wang 2017; Wang et al. 2017), video synthesis (Wang et al. 2018, 2019a), and restoration (Wang et al. 2019b; Nah et al. 2019). This broad applicability highlights its enormous technical potential and practical significance, making the intelligent processing of CV tasks a highly important area of research. The most effective solutions in this field today are based on deep learning algorithms.
@@
One of the earliest and most influential works in using deep learning for CV was LeNet (Lecun et al. 1998), which pioneered the use of CNNs for handwritten digit recognition, ushering in a new era of image processing. Following this, AlexNet’s (Krizhevsky, Sutskever, and Hinton 2017) remarkable performance on the ImageNet (Russakovsky et al. 2014) dataset not only achieved a significant leap in image classification accuracy but also cemented CNNs as the dominant method in visual processing. The subsequent introduction of ResNet (He et al. 2016) further improved image recognition accuracy and established residual-based structure as a foundational architecture for future network designs, influencing nearly all subsequent deep learning models (Ren et al. 2015; Xie et al. 2016; Szegedy et al. 2014; Gao et al. 2019; Liu et al. 2021).
@@
In recent years, the success of Transformers (Vaswani et al. 2017) in the field of natural language processing (NLP) has gradually permeated CV. Researchers have begun designing models for CV based on Transformers, such as the Vision Transformer (ViT) (Dosovitskiy et al. 2020), Swin Transformer (Liu et al. 2021) and StructViT (Kim et al. 2024), which have demonstrated capabilities comparable to CNNs in image tasks (For convenience, we collectively refer to all Transformer-based models in the CV domain as ViTs.). Currently, research in deep learning for image processing predominantly revolves around CNNs (Huang et al. 2023; Wang et al. 2022b), Transformers (Shi 2023; Wang et al. 2023), or strategies that integrate both (Lv et al. 2023; Hou et al. 2024), continuously driving technological advancements.
@@
Despite the significant advances in deep learning-based computer vision, several fundamental questions remain unanswered. For instance, why do convolutional neural networks (CNNs) require such deep architectures? What underpins the generalization ability of CNNs? Why do residual networks outperform fully convolutional networks like VGG (Simonyan and Zisserman 2014)? What are the core differences between residual-based CNNs and Transformer-based models? We believe that these questions arise from the lack of a foundational theory for CNNs.
@@
While some researchers have attempted to explain CNNs through various approaches—such as using visualization techniques to analyze the relationship between feature maps and original images (Wei et al. 2016) or exploring CNNs from the frequency domain perspective (Yin et al. 2019; Xu, Zhang, and Xiao 2019)—these explanations often have limitations and are somewhat subjective. Similarly, studies on the interpretability of Vision Transformers (ViTs) (Park and Kim 2022; Bai et al. 2022) suggest that the multi-head attention (MHA) mechanism in Transformers tends to capture low-frequency information, in contrast to CNNs, which function as high-pass filters. These studies advocate for combining the strengths of both models or improving ViTs to better capture high-frequency details. However, these methods only provide intuitive explanations for specific problems and fail to offer a theoretical framework for addressing broader issues within CNNs and ViTs. Essentially, they describe or interpret phenomena occurring within CNNs or ViTs rather than providing a comprehensive theory for these architectures.
To address these issues, we propose a unified explanation for the problems in CNNs and ViTs under the framework of the UAT. Although some previous works (Lin, Shen, and Li 2022; Yang, Feng, and Zhou 2024; Zhou 2018, 2020) have attempted to use UAT to explain certain aspects of CNNs, these studies are often limited to specific conditions and lack general applicability and practical guidance. For example, they fail to fundamentally explain the differences between residual-based CNNs and VGG, and they do not extend to ViT architectures. Our goal is to leverage the Matrix-Vector method introduced in UAT2LLMs (Wang and Li 2024b) to demonstrate that both CNNs and ViTs are concrete implementations of UAT and to theoretically resolve these issues in a way that can also provide practical guidance for network design (Wang and Li 2024a). Our contributions are as follows:
@@
  • We used the Matrix-Vector Method to demonstrate that the multi-layer convolutional and Transformer structures commonly employed in the field of CV are specific implementations of the UAT.
@@
  • We explained why CNNs require deep networks, considering the characteristics of image data and the UAT model perspective.
@@
  • We proved why the widely used residual structure in CV is so powerful.
@@
  • We compared multi-layer convolutional residual structures and Transformer structures from the UAT perspective, illustrating the source of their powerful capabilities and explaining why both models appear to have similar performance.
  • we provided a comparison between human vision and deep learning-based CV.
@@
The structure of this paper is as follows: In Section 2, we introduce the UAT and the Matrix-Vector Method, as well as their relationship. In Section 3, we use the Matrix-Vector Method to transform some commonly used operations in CV into matrix-vector forms, such as 2D convolution (3.1), 3D convolution (3.2), mean-pooling (3.3), and Transformer (3.4). In Section 4, we address some fundamental questions in CV: why CNN networks need to be deep (4.1), why residual-based CNNs are more powerful than VGG (4.2), the differences between residual-based CNNs and Transformer-based networks (4.3), and the relationship between human vision and deep learning-based CV(4.4).
2 UAT for CV
@@
We have previously introduced our goal of unifying CNNs and ViTs in the field of CV under the framework of UAT. Using UAT, we aim to explain existing problems in CV. So we will follow the UAT2LLMs’ (Wang and Li 2024b) thoughts to leverage the Matrix-Vector Method to unify CNNs and ViTs under the UAT framework. Before formally unifying CNNs and ViTs within UAT, we will briefly introduce the Matrix-Vector Method and the unification approach proposed in UAT2LLMs.
Figure 1 simply outlines the proof route in UAT2LLMs: first, $G(\mathbf{x})$ represents the general form of UAT, capable of approximating Borel measurable functions within a closed interval(Cybenko 2007; Hornik, Stinchcombe, and White 1989), where $\mathbf{I}_{n}$ denotes an n-dimensional unit cube. UAT includes one-dimensional and high-dimensional approximations, corresponding to Figures 1.a and 1.b, respectively. It is evident that the basic UAT computing unit is $\alpha_{j}\sigma(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j})$. $\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}$ can be seen as operations of convolution and mean pooling in CNNs, and MHA and linear in ViTs. Therefore, if we can transform these operations into a matrix multiplying a vector using the Matrix-Vector Method in Figures 1 (Details could be seen in UAT2LLMs), it becomes straightforward to prove that multi-layer CNNs and ViTs are specific implementations of UAT.
The Matrix-Vector Method is illustrated in Figure 1, where $T$ denotes various transformations in CNNs and ViTs, and $T_{D}$ and $T_{P}$ represent specific transformation methods. These transformations convert the input data $\mathbf{x}$ and output data $\mathbf{y}$, along with parameters $\mathbf{W}$, into vectors $\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$, and parameter matrices $\mathbf{W}^{\prime}$, satisfying the computational requirements shown in the Figure 1. To distinguish them from the original variables, we will add a symbol to the upper right of the original variables, indicating their corresponding matrix-vector form. (Note: The key challenge is converting them into matrix-vector form, so bias terms are omitted here.)
Additionally, due to the complexity of convolution and Transformer calculations, a computational tool called diamond matrix multiplication has been designed to clearly elucidate the process of converting convolution or Transformer operations into their corresponding matrix-vector representations. Its relationship with standard matrix multiplication is $\mathbf{W}\diamond\mathbf{x}=\mathbf{W}^{T}\mathbf{x}$. The diamond matrix enhances the clarity of this conversion process. (The details and properties can be found in UAT2LLMs.)
@@
3 The Matrix-Vector Method for CNNs and ViTs
In the previous sections, we explained that to unify CNNs and ViTs under the UAT, it is necessary to demonstrate that their various transformations (convolution, mean pooling, MHA, and linear) can be represented in matrix-vector form. UAT2LLMs has already shown how to represent MHA and linear in matrix-vector form. Therefore, in this section, we will specifically demonstrate how to convert convolution and mean pooling in CNNs into matrix-vector form using the Matrix-Vector Method. Additionally, we will interpret the unique aspects of MHA in ViTs within the context of CV.
Considering that convolution in CV can be either 2D or 3D and that the input and output may consist of single or multiple channels. We use the notation 1 to represent a single channel, I to represent an input with I channels, and O to represent an output with O channels. For example, 1-O indicates an operation with one channel input and $O$ channels output. Below, we will show how to convert these transformations into the corresponding matrix-vector forms.
@@ 3.2 The Matrix-Vector Method for Conv3D
According to Figure 4, the matrix-vector form of 1-O Conv3D can be easily derived and is consistent with Eq. (1). The difference lies in the fact that $\mathbf{W}_{i}$ has depth, so the corresponding $\mathbf{W}_{i}^{\prime}$ needs to be recombined as shown in Figure 4.c.
Similarly, Figures 5.a, b, and c illustrate the general process, a simple example, and the corresponding matrix-vector form of I-1 Conv3D, respectively. According to Figure 5, the matrix-vector form of I-O Conv3D is consistent with Eq. (2). The difference is that both $\mathbf{x}_{i}$ and $\mathbf{W}_{i}$ have depth, so the corresponding $\mathbf{x}_{i}^{\prime}$ and $\mathbf{W}_{i}^{\prime}$ need to be obtained as shown in Figure 5.c.
@@ 3.3 The Matrix-Vector Method for Mean Pooling
3.4 Transformer for CV
@@
In UAT2LLMs, it has been demonstrated that the feedforward network (FFN) and MHA components in the Transformer architecture can be represented in matrix-vector form. Given that ViTs are also based on the Transformer framework, we only need to focus on the differences between Transformers in CV and those in UAT2LLMs. A key feature of ViTs is that they divide the original image into multiple patches: $\mathbf{x}_{1,1},\cdots,\mathbf{x}_{n,m}$. Each image patch is then reshaped into row vectors $\bar{\mathbf{x}}_{1,1},\cdots,\bar{\mathbf{x}}_{n,m}$, which are concatenated along the column direction to obtain $\tilde{\mathbf{x}}$, as shown in Figure 7. MHA and FFN operations are then applied, so aside from the initial processing part, the subsequent transformation process completely follows the description in UAT2LLMs. Therefore, it is easy to deduce that ViTs are also specific implementations of the UAT.
3.5 Summary
We know that various modules in CV, including 2D convolution, 3D convolution, mean pooling, MHA, and linear, can all be represented by diamond matrix multiplication. Considering the relationship between diamond matrix multiplication and matrix multiplication: $\mathbf{W}\diamond\mathbf{x}=\mathbf{W}^{T}\mathbf{x}$. Therefore, unless otherwise specified, we will conveniently use $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ to represent the matrix-vector form of them. Here, $\mathbf{W}^{\prime}$ is derived from the parameters $\mathbf{W}$ of the various modules, and $\mathbf{x}^{\prime}$ is derived from the inputs $\mathbf{x}$ of the modules.

### v4

# Dynamic Universal Approximation Theory: The Basic Theory for Residual-Based Deep Learning Computer Vision Models

@@ Abstract
Computer Vision (CV) is one of the most critical branches in artificial intelligence. In recent years, various deep learning models based on convolutional architectures and Transformers have been developed to address diverse challenges in CV. These algorithms have found practical applications in fields such as robotics and facial recognition. Despite the increasing capabilities of modern CV models, several fundamental questions remain unresolved. For instance, what underpins the generalization ability of Convolutional Neural Networks (CNNs)? The root of these issues lies in the lack of a robust theoretical foundation for deep learning models in CV. To address these key challenges, we propose the Dynamic Universal Approximation Theorem (DUAT), an advance of the Universal Approximation Theorem (UAT). We demonstrate that most residual-based deep learning models in CV fall under the category of DUAT functions. By doing so, we aim to clarify some theoretical issues in CV.
@@ 1 Introduction
As a core branch of artificial intelligence, CV has a wide range of applications, encompassing tasks such as image segmentation Minaee et al. (2021); Wang et al. (2022a), classification Rawat & Wang (2017); Wang et al. (2017), video synthesis Wang et al. (2018; 2019a), and restoration Wang et al. (2019b); Nah et al. (2019). This broad applicability highlights its enormous technical potential and practical significance, making the intelligent processing of CV tasks a highly important area of research. The most effective solutions in this field today are based on deep learning algorithms.
@@
One of the earliest and most influential works in using deep learning for CV was LeNet Lecun et al. (1998), which pioneered the use of CNNs for handwritten digit recognition, ushering in a new era of image processing. Following this, AlexNet’s Krizhevsky et al. (2017) remarkable performance on the ImageNet Russakovsky et al. (2014) dataset not only achieved a significant leap in image classification accuracy but also cemented CNNs as the dominant method in visual processing. The subsequent introduction of ResNet He et al. (2016) further improved image recognition accuracy and established residual-based structure as a foundational architecture for future network designs, influencing nearly all subsequent deep learning models Ren et al. (2015); Xie et al. (2016); Szegedy et al. (2014); Gao et al. (2019); Liu et al. (2021).
@@
In recent years, the success of Transformers Vaswani et al. (2017) in the field of natural language processing (NLP) has gradually permeated CV. Researchers have begun designing models for CV based on Transformers, such as the Vision Transformer (ViT) Dosovitskiy et al. (2020), Swin Transformer Liu et al. (2021) and StructViT Kim et al. (2024), which have demonstrated capabilities comparable to CNNs in image tasks (For convenience, we collectively refer to all Transformer-based models in the CV domain as ViTs.). Currently, research in deep learning for image processing predominantly revolves around CNNs Huang et al. (2023); Wang et al. (2022b), Transformers Shi (2023); Wang et al. (2023), or strategies that integrate both Lv et al. (2023); Hou et al. (2024), continuously driving technological advancements.
@@
Despite the significant advances in deep learning-based computer vision, several fundamental questions remain unanswered. For instance, why do CNNs require such deep architectures? What underpins the generalization ability of CNNs? Why do residual networks outperform fully convolutional networks like VGG Simonyan & Zisserman (2014)? What are the core differences between residual-based CNNs and Transformer-based models? We believe that these questions arise from the lack of a foundational theory for CNNs.
@@
While some researchers have attempted to explain CNNs through various approaches—such as using visualization techniques to analyze the relationship between feature maps and original images Wei et al. (2016) or exploring CNNs from the frequency domain perspective Yin et al. (2019); Xu et al. (2019)—these explanations often have limitations and are somewhat subjective. Similarly, studies on the interpretability of ViTs Park & Kim (2022); Bai et al. (2022) suggest that the multi-head attention (MHA) mechanism in Transformers tends to capture low-frequency information, in contrast to CNNs, which function as high-pass filters. These studies advocate for combining the strengths of both models or improving ViTs to better capture high-frequency details. However, these methods only provide intuitive explanations for specific problems and fail to offer a theoretical framework for addressing broader issues within CNNs and ViTs. Essentially, they describe or interpret phenomena occurring within CNNs or ViTs rather than providing a comprehensive theory for these architectures.
To address these challenges, we propose the DUAT, which applies to both residual-based CNNs and ViTs. DUAT is an advancement of the UAT, specifically developed to address the commonly used residual structures in deep learning models for computer vision. Although previous studies Lin et al. (2022); Yang et al. (2024); Zhou (2018; 2020) have attempted to use UAT to explain certain aspects of CNNs, they are often limited by specific conditions, lacking broad applicability or practical guidance. For example, these studies have not fundamentally clarified the differences between residual-based CNNs and VGG, nor have they extended to ViTs.
Our objective is to employ the Matrix-Vector Method introduced in DUAT2LLMs Wang & Li (2024a) to demonstrate that both residual-based CNNs and ViTs fall within the class of DUAT functions, theoretically resolving the aforementioned issues and providing practical insights for network design Wang & Li (2024b). Our contributions are as follows:
  • We further developed the UAT to DUAT, which could help explain some problems in the CV area.
@@
  • We demonstrate that the residual-based multi-layer convolutional and Transformer networks commonly employed in the field of CV belong to the DUAT function.
@@
  • We explained why CNNs require deep networks, considering the characteristics of image data and the DUAT perspective.
@@
  • We proved why the widely used residual structure in CV is so powerful beyond the full convolutional networks, like VGG.
@@
  • We generally compared a residual-based CNN and a ViT from the DUAT perspective, illustrating the source of their powerful capabilities and explaining why both models appear to have similar performance.
@@
The structure of this paper is as follows: In Section 2, we introduce the UAT and the Matrix-Vector Method, as well as DUAT. In Section 3, we use the Matrix-Vector Method to transform some commonly used operations in CV into matrix-vector forms, such as 2D convolution (3.1), 3D convolution (3.2), mean-pooling (3.3), and then we give the DUAT format of residual-based CNNs (3.4) and ViTs (3.5). In Section 4, we address some fundamental questions in CV: why CNN networks need to be deep (4.1), why residual-based CNNs are more powerful than VGG (4.2) and the differences between residual-based CNNs and ViTs (4.3).
2 From UAT to DUAT
@@
Our goal is to unify residual-based CNNs and ViTs within the DUAT framework, where DUAT can be seen as an advanced form of UAT. Before introducing DUAT, we first provide an overview of UAT and briefly describe the Matrix-Vector Method proposed in DUAT2LLMs Wang & Li (2024a).
Figure 1 offers a simplified overview from DUAT2LLMs. Here, $G(\mathbf{x})$ represents the general form of UAT, which approximates Borel-measurable functions over closed intervals Cybenko (1989); Hornik et al. (1989), with $\mathbf{I}_{n}$ denoting the n-dimensional unit cube. UAT extends from one-dimensional to multi-dimensional approximations, illustrated in Figures 1.a and 1.b. The fundamental computational unit of UAT is expressed as $\alpha_{j}\sigma(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j})$. $\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}$ can be interpreted as convolution and pooling operations in CNNs, or as MHA and linear operations in ViTs, the specific mathematical forms of these basic operations differ from $\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}$. To bridge this gap, the matrix-vector approach is introduced.
The right of Figure 1 illustrates how these operations can be transformed into a matrix-vector multiplication form, $\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}$, using the Matrix-Vector Method. In this context, $T$ represents various transformations within CNNs and ViTs, while $T_{D}$ and $T_{P}$ denote specific matrix-vector transformations, varying with the choice of $T$. These transformations require converting input data $\mathbf{x}$, output data $\mathbf{y}$, and parameters $\mathbf{W}$ into their vectorized forms $\mathbf{x}^{\prime}$, $\mathbf{y}^{\prime}$, and parameter matrix $\mathbf{W}^{\prime}$. These transformations ensure the computational relation $\mathbf{W}^{\prime}\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$, as shown in Figure 1. We use a prime symbol to indicate the matrix-vector representations of these transformed variables, distinguishing them from their original forms. This establishes the link between basic network operations and $\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}$ (notably, we omit the bias term for simplification, focusing on basic operations’ transformations, like convolution).
With this connection established between network basic operations and UAT’s basic computational units, what about multi-layer networks? By expressing multi-layer networks mathematically in matrix-vector form, we observe that fully convolutional networks align with UAT in mathematical format (we illustrate this using the VGG network’s mathematical form in Section 4.2). As most current CV networks employ residual connections, writing these networks in corresponding mathematical terms reveals that their overall mathematical format aligns with UAT. However, unlike classic UAT, most parameters in these formulas vary with the input, making them functions of the input. We define this form of UAT, where some or all parameters depend on the input, as DUAT. (In Section 3.4 and 3.5, we present the DUAT mathematical form for a multi-layer residual-based CNN and ViT.) To demonstrate that a multi-layer network qualifies as DUAT, we use the matrix-vector method to represent the mathematical form of these networks, establish their consistency with UAT, and verify that certain UAT parameters vary dynamically with the input, confirming them as DUAT functions. Through the introduction of DUAT, we aim to address unresolved questions in computer vision.
@@
3 The DUAT Format of CNNs and ViTs
In the previous sections, we explained that to unify CNNs and ViTs under the DUAT, it is necessary to demonstrate that their various transformations (convolution, mean pooling, MHA, and linear) can be represented in matrix-vector form. DUAT2LLMs has already shown how to represent MHA and linear in matrix-vector form. Therefore, in this section, we will specifically demonstrate how to convert convolution and mean pooling in CNNs into matrix-vector form using the Matrix-Vector Method. Additionally, we will interpret the unique aspects of Transformer within the context of CV. In the end, we give the DUAT format of residual-based CNNs and ViTs.
Considering that convolution in CV can be either 2D or 3D and that the input and output may consist of single or multiple channels. We use the notation 1 to represent a single channel, I to represent an input with I channels, and O to represent an output with O channels. For example, 1-O indicates an operation with one channel input and $O$ channels output. Below, we will show how to convert these transformations into the corresponding matrix-vector forms. Then we will give the DUAT format of residual-based CNNs and ViTs.
@@ 3.2 The Matrix-Vector Method for Conv3D
According to Figure 4, the matrix-vector form of 1-O Conv3D can be easily derived and is consistent with Eq. 1. The difference lies in the fact that $\mathbf{W}_{i}$ has depth, so the corresponding $\mathbf{W}_{i}^{\prime}$ needs to be recombined as shown in Figure 4.c.
Similarly, Figures 5.a, b, and c illustrate the general process, a simple example, and the corresponding matrix-vector form of I-1 Conv3D, respectively. According to Figure 5, the matrix-vector form of I-O Conv3D is consistent with Eq. 2. The difference is that both $\mathbf{x}_{i}$ and $\mathbf{W}_{i}$ have depth, so the corresponding $\mathbf{x}_{i}^{\prime}$ and $\mathbf{W}_{i}^{\prime}$ need to be obtained as shown in Figure 5.c.
@@ 3.3 The Matrix-Vector Method for Mean Pooling
3.4 The DUAT Format of Residual-Based CNNs
Figure 7: A general residual module in ResNet.
Building on the previous derivation, we know that almost all types of basic operations in CNNs can be expressed in matrix-vector form. Here, we present the DUAT formulation for a general multi-layer residual-based CNN. Assuming a basic residual CNN module as shown in Figure 7, the DUAT form for a multi-layer residual-based CNN can be expressed as follows (for detailed deduction, refer to Appendix A.3):
where $i=0,1,2,\ldots$. For $i=0$, we have $\mathbf{b}^{\prime}_{1,1}=\mathbf{b}^{\prime}_{1,1}$; and for $i>0$, $\mathbf{b}^{\prime}_{j,1}$ is defined by $\mathbf{b}^{\prime}_{j,1}=\left(\mathbf{W}^{\prime}_{j,1}\mathbf{b}^{\prime}_{j-1,2}+\mathbf{b}^{\prime}_{j,1}\right)+\mathbf{W}^{\prime}_{j,1}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,2}\sigma\left(\mathbf{W}^{\prime}_{k,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,1}\right).$ Thus, $\mathbf{b}^{\prime}_{j,1}$ can be seen as a $j$-layer UAT and the input is $\mathbf{x}_{0}$. For convenience, we also consider the first term as part of the UAT. Additionally, for the term $\mathbf{b}^{\prime}_{j+1,2}$ at layer $i$, it is given by $\mathbf{b}^{\prime}_{j+1,2}=\mathbf{b}^{\prime}_{1,2}\cdots\mathbf{b}^{\prime}_{j,2}+\mathbf{b}^{\prime}_{j+1,2},j=1,2\cdots i$. Eq. 3 aligns with the UAT form, as illustrated in Figure 1. In this context, $\mathbf{b}_{j,1}$ functions as a dynamic parameter that is approximated through UAT. Consequently, Eq. 3 satisfies the DUAT criterion, adopting a UAT form where the parameters dynamically adapt to the input.
3.5 The DUAT Format of ViTs
In DUAT2LLMs, it has been demonstrated that the feedforward network (FFN) and MHA components in the Transformer architecture can be represented in matrix-vector form. Given that ViTs are also based on the Transformer framework, we only need to focus on the differences between Transformers in CV and those in LLMs. A key feature of ViTs is that they divide the original image into multiple patches: $\mathbf{x}_{1,1},\cdots,\mathbf{x}_{n,m}$. Each image patch is then reshaped into row vectors $\bar{\mathbf{x}}_{1,1},\cdots,\bar{\mathbf{x}}_{n,m}$, which are concatenated along the column direction to obtain $\tilde{\mathbf{x}}$, as shown in Figure 8. MHA and FFN operations are then applied, so aside from the initial processing part, the subsequent transformation process completely follows the description in DUAT2LLMs.
@@ 3.5 The DUAT Format of ViTs
Figure 9: The process of Transformer.
Therefore, a multi-layer ViT structured in Figure 9 can be expressed as:
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,3}^{\prime})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (4)
When $i=0$, the parameters are defined as follows: $\mathbf{W}^{\prime}_{1,1}=\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{1,3}=\mathbf{b}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,2}=\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{1,2}=\mathbf{b}^{\prime}_{1,2}$. We define the parameters for cases where $i\geq 1$. For each $j=1,2,\ldots,i$, the updates are as follows: $\mathbf{W}^{\prime}_{j+1,1}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}_{j,1}$, $\mathbf{b}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{b}_{j,3}+\mathbf{b}^{\prime}_{j+1,3}$, $\mathbf{W}^{\prime}_{j+1,2}=\mathbf{W}^{\prime}_{j+1,2}\mathbf{W}_{j,1}$, and $\mathbf{W}^{\prime}_{j,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}^{\prime}_{j,3}$. Additionally, for $j=2,\ldots,i+1$, the bias terms $\mathbf{b}^{\prime}_{j,2}$ are updated according to:
$\mathbf{b}^{\prime}_{j,2}=$ $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})$ (5)
$+$ $\mathbf{W}^{\prime}_{j,2}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2}).$
It means that the whole computing process is serial, if we want to compute $\mathbf{W}^{\prime}_{3,1}$, we must calculate $\mathbf{W}^{\prime}_{2,1}$, because $\mathbf{W}^{\prime}_{3,1}=\mathbf{W}^{\prime}_{3,1}\mathbf{W}^{\prime}_{2,1}$. Thus, we have established that a multilayer ViT is also a DUAT function. The mathematical representation of the ViT aligns with the DUAT framework, with $\mathbf{b}^{\prime}_{j,2}$ closely approximated by DUAT. The most notable difference between ViTs and residual-based CNNs is the dynamic input-dependent nature of the parameters in the MHA mechanism. Consequently, in the DUAT formulation associated with ViT, the parameters $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ for $i+1\geq j\geq 1$, and $\mathbf{W}^{\prime}_{j,3}$ for $i+1>j>1$ in layer $i$, all adapt dynamically based on the input. For further details, please refer to Appendix A.4.